# Stuart-Landau Oscillatory Graph Neural Network

### v1

@@ 5.3. Graph-level tasks
To further evaluate performance of SLGNNs on graph-level tasks, We benchmark graph classification performance on the MoleculeNet datasets (wuMoleculeNetBenchmarkMolecular2017) (MUTAG, ENZYMES, PROTEINS), and graph regression on datasets from the Open Graph Benchmark (ivanovUnderstandingIsomorphismBias2019; huOpenGraphBenchmark2020) (ESOL, FreeSolv, Lipophilicity). For these experiments, we adopt a standard 80:10:10 train-validation-test split with a batch size of 64. As presented in Table 4 and 5, the results demonstrate the strong performance of SLGNNs on these tasks, which consistently outperform other OGNNs by up to 80% (see ENZYMES-Tran cells).
@@
Table 4. Graph classification accuracy (%) on MoleculeNet (wuMoleculeNetBenchmarkMolecular2017) datasets. The means and standard deviations are averaged over 50 random splits and weight initializations.
@@
Table 5. Graph regression (L2 loss) on OGBG-MOL datasets (ivanovUnderstandingIsomorphismBias2019; huOpenGraphBenchmark2020). The means and standard deviations are averaged over 50 random splits and weight initializations.
@@ 6. Discussions
Beyond depth scalability, we also evaluate model robustness against graph perturbations by randomly adding fake edges to the graph structure, following the protocol from fengGraphRandomNeural2020 (fengGraphRandomNeural2020). As shown in Figure 2, SLGNN exhibits superior robustness compared to Kuramoto and achieves a level of robustness comparable to that of GraphCON.
@@
A key property of the Stuart Landau Oscillator is its ability to operate near a critical state, which has been extensively used by the the neuroscience community for modeling the brain (deco). At this critical point, the system’s dynamics converge at a slower, algebraic rate rather than an exponential one. This state is reached when the following condition is met: for each node $j$,
@@
Directly verifying this condition is challenging, as the coupling strength $\kappa$ is implicitly set by the magnitude of the learned weights for SL-GCN and -GAT, and $A_{jk}$ is dependent on the attention mechanism for SL-GAT and -Tran, both of which may not be fixed across layers. However, we find compelling empirical evidence that the model can leverage this property and reach a near-critical state. For instance, in our experiments with SL-GCN on Cora, Citeseer, and Pubmed, the hyperparameter tuner consistently selects a Hopf-parameter $\alpha$ close to 0 (e.g. 0.04), while we concurrently observe the average absolute value of $cos(\phi_{j}-\phi_{k})\frac{r_{k}}{r_{j}}$ remains around 1.0 (e.g. 1.06). In this special case, the oscillators appear to be almost synchronized, but with $\alpha\sim 0$, the Stuart-Landau oscillator is near its critical state. This suggests the hyperparameter tuner has leveraged the slower algebraic convergence to improve the performance of the model, a characteristic that no other OGNN frameworks have. This type of criticality-seeking behaviour appears to align with the idea that the brain operates near criticality (RevModPhys.90.031001). The precise implications in deep learning will be the topic of a future study.
@@ 8. Acknowledgement
KZ was supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot-Watt University and the University of Edinburgh. DNR was funded by National Science Centre, Poland, grant number 2023/50/A/ST1/00447 and partially supported by the Modeling Nature (MNat) Research Unit, project QUAL21-011. PD and FT are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). PD was supported by the INdAM-GNCS project “NLA4ML—Numerical Linear Algebra Techniques for Machine Learning”. This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh, as well as AI Research Resource Programme (AIRR) under UK Research and Innovation (UKRI).
@@ Appendix A Hyperparameter Tuning Setup
We conduct extensive hyperparameter search across all model-dataset pairs presented in Table 2-5 with Ray platform (moritzRayDistributedFramework2017) and Optuna (akibaOptunaNextgenerationHyperparameter2019) search algorithm over 1000 trials for node-level tasks and 500 for graph-level ones. The search spaces for all GCN, GAT, and Tran variants include: latent dimension size [16,128], number of layers [1,80], learning rate [10-5, 10-2], weight decay [10-6, 1.0], LeakyReLU gradient [0.0, 1.0], dropout and input dropout rates [0.0, 0.5]. The time step size is fixed at $dt=$ 1.0. For GAT and Tran, we fix the number of attention heads to 2 or 4. For Tran, we additionally tune the attention dimension between [4, 80] and the coupling strength $\kappa$ in [0.1,1.5]. For the OGNNs, we tuned their specific parameters:
@@
  • For GraphCON, we tune both $\alpha$ and $\gamma$ in [0.0, 1.5], consistent with the search space used in (ruschGraphCoupledOscillatorNetworks2022).
@@ Appendix C Additional Results on Model Depths
We note that SLGNN’s optimal configurations are less concentrated in the shallow 1- to 10-layer range compared to GraphCON, Kuramoto, and the Baseline. This reinforces our claim that SLGNNs typically leverage deeper network structures to achieve better performance. On the other hand, there are a significant number of Kuramoto’s configurations in the very deep 70- to 80-layer range. At first glance, this may seem to contradict the conclusions drawn from Figure 1. However, we observe that models with optimal depths in this range are predominantly those using Tran coupling, a configuration that makes models more robust at greater depths comparing to the GCN or GAT coupling. It is also known that certain datasets, such as the heterophilic ones, are less susceptible to performance degradation in deep architectures (guoTamingOversmoothingRepresentation2023). Most importantly, increased depth does not always translate to better performance, particularly for the Baseline and other OGNN models.

### v2

@@ 5.3. Graph-level tasks
To further evaluate performance of SLGNNs on graph-level tasks, We benchmark graph classification performance on the MoleculeNet datasets (Wu et al., 2017) (MUTAG, ENZYMES, PROTEINS), and graph regression on datasets from the Open Graph Benchmark (Ivanov et al., 2019; Hu et al., 2020) (ESOL, FreeSolv, Lipophilicity). For these experiments, we adopt a standard 80:10:10 train-validation-test split with a batch size of 64. As presented in Table 4 and 5, the results demonstrate the strong performance of SLGNNs on these tasks, which consistently outperform other OGNNs by up to 80% (see ENZYMES-Tran cells).
@@
Table 4. Graph classification accuracy (%) on MoleculeNet (Wu et al., 2017) datasets. The means and standard deviations are averaged over 50 random splits and weight initializations.
@@
Table 5. Graph regression (L2 loss) on OGBG-MOL datasets (Ivanov et al., 2019; Hu et al., 2020). The means and standard deviations are averaged over 50 random splits and weight initializations.
@@ 6. Discussions
Beyond depth scalability, we also evaluate model robustness against graph perturbations by randomly adding fake edges to the graph structure, following the protocol from Feng et al. (Feng et al., 2020). As shown in Figure 2, SLGNN exhibits superior robustness compared to Kuramoto and achieves a level of robustness comparable to that of GraphCON.
@@
A key property of the Stuart Landau Oscillator is its ability to operate near a critical state, which has been extensively used by the the neuroscience community for modeling the brain (Ponce-Alvarez and Deco, 2024). At this critical point, the system’s dynamics converge at a slower, algebraic rate rather than an exponential one. This state is reached when the following condition is met: for each node $j$,
@@
Directly verifying this condition is challenging, as the coupling strength $\kappa$ is implicitly set by the magnitude of the learned weights for SL-GCN and -GAT, and $A_{jk}$ is dependent on the attention mechanism for SL-GAT and -Tran, both of which may not be fixed across layers. However, we find compelling empirical evidence that the model can leverage this property and reach a near-critical state. For instance, in our experiments with SL-GCN on Cora, Citeseer, and Pubmed, the hyperparameter tuner consistently selects a Hopf-parameter $\alpha$ close to 0 (e.g. 0.04), while we concurrently observe the average absolute value of $cos(\phi_{j}-\phi_{k})\frac{r_{k}}{r_{j}}$ remains around 1.0 (e.g. 1.06). In this special case, the oscillators appear to be almost synchronized, but with $\alpha\sim 0$, the Stuart-Landau oscillator is near its critical state. This suggests the hyperparameter tuner has leveraged the slower algebraic convergence to improve the performance of the model, a characteristic that no other OGNN frameworks have. This type of criticality-seeking behaviour appears to align with the idea that the brain operates near criticality (Muñoz, 2018). The precise implications in deep learning will be the topic of a future study.
@@ 8. Acknowledgement
KZ was supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot-Watt University and the University of Edinburgh. DNR was funded by National Science Centre, Poland, grant number 2023/50/A/ST1/00447 and partially supported by the Modeling Nature (MNat) Research Unit, project QUAL21-011. PD and FT are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). FT is partially funded by the PRIN-MUR project MOLE (code 2022ZK5ME7) and the PRIN-PNRR project FIN4GEO (code P2022BNB97). PD was supported by the INdAM-GNCS project “NLA4ML—Numerical Linear Algebra Techniques for Machine Learning”. This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh, as well as AI Research Resource Programme (AIRR) under UK Research and Innovation (UKRI).
@@ Appendix A Hyperparameter Tuning Setup
We conduct extensive hyperparameter search across all model-dataset pairs presented in Table 2-5 with Ray platform (Moritz et al., 2017) and Optuna (Akiba et al., 2019) search algorithm over 1000 trials for node-level tasks and 500 for graph-level ones. The search spaces for all GCN, GAT, and Tran variants include: latent dimension size [16,128], number of layers [1,80], learning rate [10-5, 10-2], weight decay [10-6, 1.0], LeakyReLU gradient [0.0, 1.0], dropout and input dropout rates [0.0, 0.5]. The time step size is fixed at $dt=$ 1.0. For GAT and Tran, we fix the number of attention heads to 2 or 4. For Tran, we additionally tune the attention dimension between [4, 80] and the coupling strength $\kappa$ in [0.1,1.5]. For the OGNNs, we tuned their specific parameters:
@@
  • For GraphCON, we tune both $\alpha$ and $\gamma$ in [0.0, 1.5], consistent with the search space used in (Rusch et al., 2022).
@@ Appendix C Additional Results on Model Depths
We note that SLGNN’s optimal configurations are less concentrated in the shallow 1- to 10-layer range compared to GraphCON, Kuramoto, and the Baseline. This reinforces our claim that SLGNNs typically leverage deeper network structures to achieve better performance. On the other hand, there are a significant number of Kuramoto’s configurations in the very deep 70- to 80-layer range. At first glance, this may seem to contradict the conclusions drawn from Figure 1. However, we observe that models with optimal depths in this range are predominantly those using Tran coupling, a configuration that makes models more robust at greater depths comparing to the GCN or GAT coupling. It is also known that certain datasets, such as the heterophilic ones, are less susceptible to performance degradation in deep architectures (Guo et al., 2023). Most importantly, increased depth does not always translate to better performance, particularly for the Baseline and other OGNN models.
@@ Appendix D Implicit Solving Magnitude Equation
Appendix E Computational Complexity
The computational complexity of the proposed SLGNN is consisted of the coupling part and the IMEX part. The coupling complexity depends on the specific coupling function used. For instance, a GCN coupling scales as $\mathcal{O}(|E|d+|V|d^{2})$, where $|V|$ and $|E|$ denote the number of nodes and edges, respectively, and $d$ represents the feature dimension.
Regarding the IMEX scheme, the implicit step using Newton’s method has a per-feature, per-iteration complexity of $\mathcal{O}(1)$. Consequently, the complexity of the IMEX scheme over the entire graph is $\mathcal{O}(k|V|d)$, where $k$ represents the number of Newton iterations. Our empirical observations suggest that $k$ depends primarily on the feature distribution rather than the graph size.
Empirically, SLGNN is observed to be up to $2\times$ slower per epoch than Kuramoto and up to $3\times$ slower than GraphCON for the best hyperparameter configurations as shown in Table 8. However, it remains substantially faster than the explicit high-order ODE solvers used in prior works, such as the Dormand Prince method. In terms of memory efficiency, the VRAM usage of SLGNN is comparable to Kuramoto and within $1.5\times$ of GraphCON as shown in Table 9. It is important to note that SLGNN often leverages deeper architectures to achieve optimal performance as shown in Figure 3, which naturally contributes to increased computation time. This additional computational cost is counterbalanced by the model’s superior predictive performance.
Table 8. Runtime per epoch (ms) comparison. Values represent mean $\pm$ standard deviation.
Model Cora Citeseer Pubmed
Baseline-GCN $9.62\pm 0.74$ $11.10\pm 0.38$ $18.39\pm 0.55$
GraphCON-GCN $11.94\pm 0.51$ $13.68\pm 1.13$ $20.42\pm 0.56$
Kuramoto-GCN $16.22\pm 0.48$ $19.00\pm 0.31$ $24.73\pm 0.41$
SL-GCN $32.82\pm 2.32$ $32.40\pm 1.10$ $41.22\pm 1.79$
Model Texas Wisconsin Cornell
Baseline-GCN $0.40\pm 0.72 11$$59\pm 0.53 11.$$1.16\pm 0.49$
GraphCON-GCN $12.92\pm 0.57$ $13.73\pm 0.24$ $13.43\pm 0.57$
Kuramoto-GCN $18.67\pm 0.81$ $18.76\pm 0.40$ $18.95\pm 0.70$
SL-GCN $37.59\pm 1.89$ $36.98\pm 1.20$ $39.85\pm 1.85$
Table 9. Peak VRAM Usage (MB) comparison. Values represent mean $\pm$ standard deviation.
Model Cora Citeseer Pubmed
Baseline-GCN $66.42\pm 0.00$ $98.50\pm 0.00$ $262.79\pm 0.00$
GraphCON-GCN $68.26\pm 0.00$ $97.53\pm 0.00$ $278.06\pm 0.00$
Kuramoto-GCN $129.34\pm 0.00$ $259.35\pm 0.00$ $390.16\pm 0.00$
SL-GCN $123.06\pm 2.74$ $179.64\pm 5.61$ $574.39\pm 10.20$
Model Texas Wisconsin Cornell
Baseline-GCN $22.92\pm 0.00$ $26.75\pm 0.00$ $27.14\pm 0.00$
GraphCON-GCN $22.42\pm 0.00$ $26.19\pm 0.00$ $27.31\pm 0.00$
Kuramoto-GCN $27.41\pm 0.00$ $33.21\pm 0.00$ $32.30\pm 0.00$
SL-GCN $31.32\pm 2.14$ $36.12\pm 2.55$ $34.75\pm 1.41$